OpenStudy (anonymous):

Write the equation in a standard form ax^2+bx+c=0 , then give the values of a,b, and c. 3x-1/x+5 = x-3/x+4

OpenStudy (lalaly):

\[\large{\frac{3x-1}{x+5}}=\frac{x-3}{x+4}\] \[\large{(3x-1)(x+4)=(x-3)(x+5)}\]

OpenStudy (lalaly):

\[\large{3x^2+12x-x-4}=x^2+5x-3x-15\]

OpenStudy (lalaly):

\[\large{3x^2+11x-4}=x^2+2x-15\]

OpenStudy (lalaly):

\[\large{2x^2+9x+11}=0\]

OpenStudy (anonymous):

i got it too..